OpenStudy (anonymous):

Given eight students, three of which are females, if two students are selected at random, WITHOUT REPLACEMENT, what is the probability that both students are female?

OpenStudy (anonymous):

Do you want this as a percentage?

OpenStudy (anonymous):

Fraction and decimal

OpenStudy (anonymous):

3/8 * 2/7

OpenStudy (anonymous):

3/8* 2/7 =3/28

OpenStudy (anonymous):

3/28= 0.107